# Enhancing Photometric Redshift Catalogs Through Color-Space Analysis: Application to KiDS-Bright Galaxies

### v1

@@ Appendix A Removing 20 per cent of the KiDS-Bright sample
In Sec. 4.2 we presented a method to clean up the KiDS-Bright sample for three main criteria, based on magnitude limit and photo-$z$ performance, and on their combination. In each case, we adjusted the cuts in such a way as to remove roughly 10% of the galaxies from the input dataset. Here we present the results of the same procedure, but for the case when we remove about 20% objects from KiDS-Bright. The outcome is illustrated in Figs. 12-15 and summarized in Table 3, which are counterparts of those shown in Sec. 4.2 for the $\sim 10\%$ removal. We illustrate the SOM distribution of the removed cells in Fig. 16 for a combination of 20% galaxies removed. As can be seen, even such a more aggressive cut of keeping $\sim 80\%$ of galaxies does not lead to an enormous improvement in photo-$z$ quality in terms of scatter (SMAD of $\delta z/(1+z)$), which is never reduced by more than 5% despite removing 1/5th of the KiDS-Bright galaxies. The completeness ratio shown in Fig. 17, also shows the negligible improvement of the completeness ratio in stellar mass and photo-$z$ as compared to Fig. 4.
@@
Combined (GAMA)1 $8.07\times 10^{5}$ 0.237 $-1.6\times 10^{-4}$ $3.3\times 10^{-4}$ 0.0215 0.0172
Combined (Espec)2 $8.11\times 10^{5}$ 0.234 $-1.0\times 10^{-4}$ $3.8\times 10^{-4}$ 0.0214 0.0172
@@
No cut $5.87~{}[4.13]\times 10^{5}$ 0.209 [0.288] $1.7~{}[13.3]\times 10^{-4}$ $6.0~{}[17.6]\times 10^{-4}$ 0.0216 [0.0273] 0.0166 [0.0216]
$\langle$ r $\rangle$<⟩ <19.75 $5.81~{}[2.16]\times 10^{5}$ 0.208 [0.260] $1.5~{}[16.1]\times 10^{-4}$ $5.8~{}[19.9]\times 10^{-4}$ 0.0214 [0.0270] 0.0166 [0.0211]
$|Dz|$ (GAMA) $<$3.93% $4.81~{}[3.17]\times 10^{5}$ 0.221 [0.307] $-6.8~{}[7.8]\times 10^{-4}$ $-1.4~{}[12.7]\times 10^{-4}$ 0.0202 [0.0254] 0.0163 [0.0208]
$|Dz|$ (Espec) $<$3.93% $4.87~{}[3.15]\times 10^{5}$ 0.221 [0.296] $-6.1~{}[8.0]\times 10^{-4}$ $-0.9~{}[12.6]\times 10^{-4}$ 0.0198 [0.0252] 0.0160 [0.0206]
$\Delta$ z (GAMA) $<$0.006 $5.08~{}[2.91]\times 10^{5}$ 0.210 [0.290] $-5.4~{}[5.2]\times 10^{-4}$ $-0.1~{}[10.5]\times 10^{-4}$ 0.0204 [0.0252] 0.0165 [0.0208]
$\Delta$ z (Espec) $<$0.007 $5.15~{}[2.85]\times 10^{5}$ 0.206 [0.274] $-3.8~{}[3.9]\times 10^{-4}$ $1.0~{}[9.3]\times 10^{-4}$ 0.0201 [0.0250] 0.0162 [0.0206]
Combined (GAMA)1 $5.42~{}[2.65]\times 10^{5}$ 0.214 [0.283] $-4.6~{}[8.4]\times 10^{-4}$ $0.4~{}[12.9]\times 10^{-4}$ 0.0202 [0.0251] 0.0163 [0.0207]
Combined (Espec)2 $5.39~{}[2.72]\times 10^{5}$ 0.214 [0.274] $-4.0~{}[8.7]\times 10^{-4}$ $0.9~{}[13.2]\times 10^{-4}$ 0.0201 [0.0252] 0.0162 [0.0208]
@@
Figure 12: Same as Fig. 8 but the black [red] histogram shows the galaxies from the cells with the mean $r$ magnitude to be less [more] than 19.75, removing $\sim 20\%$ galaxies.
Figure 13: Same as Fig. 9 but removing 20% KiDS Bright galaxies. Left panels: The black [red] histogram shows the galaxies from the cells with the mean relative difference between photometric and spectroscopic redshift from GAMA to be less [more] than 3.93%. Right panels: Same as the left panel, but the spectroscopic redshift is from the Espec sample.
Figure 14: Same as Fig. 10 but removing 20% galaxies with the mean absolute difference between photometric and spectroscopic redshift from GAMA to be less [more] than 0.006. The right panel shows the same but the mean absolute difference between photometric and spectroscopic redshift from Espec is less [more] than 0.007.
Figure 15: Same as Fig. 11 but instead of 10% removal of galaxies, removing $\sim$19galaxies. Left panels: The tan histograms show the distribution of KiDS galaxies versus r-magnitude, g-r color, stellar mass, and photometric redshifts. The red histogram shows the galaxies from the cells with either $\langle r\rangle>19.84$ (green), $|Dz|>6.48\%$ (purple), and based on $|\Delta z|>0.01$ (blue), removing $\sim 19\%$ galaxies. The 80% of the clean samples are shown in black. Right panels: Same as the left panel, but the spectroscopic redshift is the mean from the Especz sample instead of GAMA in each SOM cell with $|Dz|>6.48\%$ (purple), and based on $|\Delta z|>0.011$, removing $\sim 19\%$ galaxies.
@@ Appendix B Dependence of photo-$z$ quality on the percentage of removed galaxies
Here we generalize the possible sample cleanup and check how many galaxies one should remove from KiDS-Bright to obtain the desired reduction in photo-$z$ bias or scatter. This is summarized in Fig. 18, where we show the SMAD and mean photo-$z$ residual (both based on GAMA) for various percentages of removed galaxies, starting from 5% up to as much as 50%, that is reducing the dataset by half. As the left panel indicates, we find that SMAD values are consistently lowered if we remove the faint end, and this reduction is up to 10% but only when the faintest half of the sample is removed ($\langle r\rangle>19.36$ mag). On the other hand, cuts $\Delta z$ or $Dz$ hardly affect the SMAD which changes by less than 3% irrespective of the culled fraction. The picture is reversed if we consider the change in mean residuals. In the right panel of Fig. 18 we can see that $\langle\Delta z\rangle$ stays at the level of $\sim 8\times 10^{-4}$ no matter how many faint galaxies are removed. On the other hand, this photo-$z$ residual goes down roughly linearly with more aggressive cuts in SOM-based $Dz$ and $\Delta z$. In fact, at around 40 - 45% of removed galaxies, it crosses the zero line and becomes negative.
Figure 18: Left panel: Relative change in the SMAD of $\delta z/(1+z)$ of the KiDS-Bright sample as a function of the fraction of removed galaxies, after applying cuts (based on GAMA) on the photometric dataset, each employing SOM cell averages, based on $\langle r\rangle$ (blue-circle), relative difference $|Dz|$ (orange-square) and absolute $|\Delta z|$ (green-diamond). Right panel: Variation of the mean photo-$z$ residual for the same setups as in the left panel.

### v2

@@ Appendix A Removing 20 percent of the KiDS-Bright sample
In Sect. 4.2, we present a method for cleaning up the KiDS-Bright sample for three main criteria, based on magnitude limit and photo-$z$ performance, and on their combination. In each case, we adjust the cuts in such a way as to remove roughly 10% of the galaxies from the input dataset. Here we present the results of the same procedure, but for the case when we remove about 20% objects from KiDS-Bright. The outcome is illustrated in Figs. 12-15 and summarize in Table 3, which are counterparts of those shown in Sect. 4.2 for the $\sim 10\%$ removal. We illustrate the SOM distribution of the removed cells in Fig. 16 for a combination of 20% galaxies removed. As can be seen, even such a more aggressive cut of keeping $\sim 80\%$ of galaxies does not lead to an enormous improvement in photo-$z$ quality in terms of scatter (SMAD of $\delta z/(1+z)$), which is never reduced by more than 5% despite removing 1/5th of the KiDS-Bright galaxies. The completeness ratio shown in Fig. 17 also shows the negligible improvement of the completeness ratio in stellar mass and photo-$z$ as compared to Fig. 4.
@@
Figure 12: Same as Fig. 8, but here the black [red] histogram shows the galaxies from the cells with the mean $r$ magnitude to be less [more] than 19.75, removing $\sim 20\%$ galaxies.
Figure 13: Same as Fig. 9, but with 20% of the KiDS Bright galaxies removed. Upper panels: Galaxies from the cells with the mean relative difference between photometric and spectroscopic redshift from GAMA to be less [more] than 3.93%, shown as a black [red] histogram. Bottom panels: Same as the upper panel, but the spectroscopic redshift is from the Espec sample.
Figure 14: Same as Fig. 10, but removing 20% of galaxies with the mean absolute difference between photometric and spectroscopic redshift from GAMA to be less [more] than 0.006. The bottom panel shows the same but the mean absolute difference between photometric and spectroscopic redshift from Espec is less [more] than 0.007.
Figure 15: Same as Fig. 11 but instead of 10% removal of galaxies, removing $\sim$19galaxies. Upper panels: The tan histograms show the distribution of KiDS galaxies versus r-magnitude, g-r color, stellar mass, and photometric redshifts. The red histogram shows the galaxies from the cells with either $\langle r\rangle>19.84$ (green), $|Dz|>6.48\%$ (purple), and based on $|\Delta z|>0.01$ (blue), removing $\sim 19\%$ galaxies. The 80% of the clean samples are shown in black. Bottom panels: Same as the upper panel, but the spectroscopic redshift is the mean from the Espec sample instead of GAMA in each SOM cell with $|Dz|>6.48\%$ (purple), and based on $|\Delta z|>0.011$, removing $\sim 19\%$ galaxies.
@@
Combined (GAMA-based)1 $8.07\times 10^{5}$ 0.237 $-1.6\times 10^{-4}$ $3.3\times 10^{-4}$ 0.0215 0.0172
Combined (Espec-based)2 $8.11\times 10^{5}$ 0.234 $-1.0\times 10^{-4}$ $3.8\times 10^{-4}$ 0.0214 0.0172
@@
(×$\times 10^{5}$) (×$\times 10^{-4}$) (×$\times 10^{-4}$)
No cut $5.87\leavevmode\nobreak\ [4.13]$ 0.209 [0.288] $1.7\leavevmode\nobreak\ [13.3]$ $6.0\leavevmode\nobreak\ [17.6]$ 0.0216 [0.0273] 0.0166 [0.0216]
$\langle$ r $\rangle$<⟩ <19.75 $5.81\leavevmode\nobreak\ [2.16]$ 0.208 [0.260] $1.5\leavevmode\nobreak\ [16.1]$ $5.8\leavevmode\nobreak\ [19.9]$ 0.0214 [0.0270] 0.0166 [0.0211]
$|Dz|$ $<$3.93% (GAMA-based) $4.81\leavevmode\nobreak\ [3.17]$ 0.221 [0.307] $-6.8\leavevmode\nobreak\ [7.8]$ $-1.4\leavevmode\nobreak\ [12.7]$ 0.0202 [0.0254] 0.0163 [0.0208]
$|Dz|$ $<$3.93% (Espec-based) $4.87\leavevmode\nobreak\ [3.15]$ 0.221 [0.296] $-6.1\leavevmode\nobreak\ [8.0]$ $-0.9\leavevmode\nobreak\ [12.6]$ 0.0198 [0.0252] 0.0160 [0.0206]
$\Delta z$ $<$0.006 (GAMA-based) $5.08\leavevmode\nobreak\ [2.91]$ 0.210 [0.290] $-5.4\leavevmode\nobreak\ [5.2]$ $-0.1\leavevmode\nobreak\ [10.5]$ 0.0204 [0.0252] 0.0165 [0.0208]
$\Delta z$ $<$0.007 (Espec-based) $5.15\leavevmode\nobreak\ [2.85]$ 0.206 [0.274] $-3.8\leavevmode\nobreak\ [3.9]$ $1.0\leavevmode\nobreak\ [9.3]$ 0.0201 [0.0250] 0.0162 [0.0206]
Combined (GAMA-based)1 $5.42\leavevmode\nobreak\ [2.65]$ 0.214 [0.283] $-4.6\leavevmode\nobreak\ [8.4]$ $0.4\leavevmode\nobreak\ [12.9]$ 0.0202 [0.0251] 0.0163 [0.0207]
Combined (Espec-based)2 $5.39\leavevmode\nobreak\ [2.72]$ 0.214 [0.274] $-4.0\leavevmode\nobreak\ [8.7]$ $0.9\leavevmode\nobreak\ [13.2]$ 0.0201 [0.0252] 0.0162 [0.0208]
@@ Appendix B Dependence of photo-$z$ quality on the percentage of removed galaxies
Here, we generalize the possible sample cleanup and check how many galaxies we should remove from KiDS-Bright to obtain the desired reduction in photo-$z$ bias or scatter. This is summarized in Fig. 18, where we show the SMAD and mean photo-$z$ residual (both based on GAMA) for various percentages of removed galaxies, starting from 5% up to as much as 50%, which ends up reducing the dataset by half. As the left panel indicates, we find that SMAD values are consistently lowered if we remove the faint end, and this reduction is up to 10% but only when the faintest half of the sample is removed ($\langle r\rangle>19.36$ mag). On the other hand, cuts $\Delta z$ or $Dz$ hardly affect the SMAD which changes by less than 3% irrespective of the culled fraction. The picture is reversed if we consider the change in mean residuals. In the right panel of Fig. 18 we can see that $\langle\Delta z\rangle$ stays at the level of $\sim 8\times 10^{-4}$ no matter how many faint galaxies are removed. On the other hand, this photo-$z$ residual goes down roughly linearly with more aggressive cuts in SOM-based $Dz$ and $\Delta z$. In fact, at around 40 - 45% of removed galaxies, it crosses the zero line and becomes negative.
Figure 18: Relative change in the SMAD of $\delta z/(1+z)$ of the KiDS-Bright sample as a function of the fraction of removed galaxies (left), after applying cuts (based on GAMA) on the photometric dataset, each employing SOM cell averages, based on $\langle r\rangle$ (blue-circle), relative difference $|Dz|$ (orange-square) and absolute $|\Delta z|$ (green-diamond). Right panel: Variation in the mean photo-$z$ residual for the same setups as in the left panel.